OpenStudy (anonymous):

Find the area of a rhombus with diagonals 9 ft. and 12 ft.

OpenStudy (anonymous):

\[A = \frac12d_{1}d_{2}\]

OpenStudy (anonymous):

A = 1/2(9)(12) A = 9(6) A = 54